log2(x+1) - log2(x-1) = 4
Answer by edjones(7311)

log2(x+1) - log2(x-1) = 4
2^(log2(x+1)) / 2^( log2(x-1)) = 2^4
(x+1)/(x-1)=16
x+1=16(x-1)
x+1=16x-16
-15x=-17
x=17/15
.
Ed